OpenStudy (anonymous):

King’s Service Station uses two kinds of cleaning solution containing acid and water. “Attack” is 2% acid and “Blast” is 6% acid. They want to mix the two to get 60 qt of a solution that is 5% acid. How many quarts of each should they use?

OpenStudy (asnaseer):

lets say we used 'x' quarts of the "Attack" solution. that means we used (60-x) quarts of the "Blast" solution. so the concentration of acid can be written as:\[\frac{2}{100}*x+\frac{6}{100}*(60-x)=\frac{5}{100}*60\]multiply both sides by 100 to simplify this to:\[2x+6(60-x)=5*60\]now we can simplify further by dividing both sides by 2 to get:\[x+3(60-x)=5*30\]\[x+180-3x=150\]\[180-2x=150\]\[2x=180-150=30\]\[x=15\] so we need 15 quarts of "Attack" and (60-15)=45 quarts of "Blast"